Circular No. 8202

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS R. M. Wagner, LBT Observatory; G. Schwarz, University of Arizona;

S. Starrfield, Arizona State University; U. Munari, E. Giro, and A. Siviero, Asiago Observatory; A. Henden, USRA and U.S. Naval Observatory, Flagstaff (USNOFS); P. Szkody, University of Washington; and H. Bond, Space Telescope Science Institute, write that optical spectra, obtained with the 6.5-m MMT (range 383–905 nm, resolution 0.6 nm) on Sept. 13.49 UT and with the Asiago 1.82-m telescope (+ AFOSC) on Sept. 12.13 (range 485–780 nm, resolution 0.8 nm) and Sept. 13.15 (in échelle mode; range 360–860 nm; resolving power 3800), reveal a dramatic change when compared to their previous spectra (cf. IAUC 7982, 7992, 8005). In addition to the Balmer and He I absorption lines characteristic of a B3 V star (shortward of 600 nm) and the TiO and VO bands (longward of 600 nm) that dominated their previous spectra, the spectrum now exhibits absorption bands of TiO, VO, ZrO, YO, and perhaps CrO, LaO, and ScO in the region 490–600 nm that were previously absent. Overlapping bands make positive identification difficult. The new spectra suggest that, while the B3 V star is still visible in the U and B bands, the O-rich and very cool atmosphere of the outbursting star is now also dominating the V band (and not only the R and I bands, as it has been since Oct. 2002). The changes are accompanied by a 0.5-mag brightening in V, R, and I bands since last May (on Sept. 11–13, V = 15.50, B −V = +1.00, V −R_{c} = +2.70, R_{c} −I_{c} = +3.25; formal errors 0.01 mag; USNOFS 1.0-m telescope), but with no appreciable changes in the near-infrared (on Sept. 13.51, J = 6.76, H = 5.84, K′ = 5.10, L′ = 3.54; errors 0.03 mag; USNOFS 1.55-m telescope + ASTROCAM). Monitoring of the changes in progress is urged.

IGR J17544–2619 R. González-Riestra, M. Santos-Lleó, L. Tomás, and R. Pérez-Martín-

ez, XMM-Newton Science Operations Centre, Madrid, report that IGR J17544–2619 was clearly detected (countrate 0.41 counts/s; EPIC-pn + thin filter) in the field-of-view of V4643 Sgr during a 10000-s XMM-Newton observation on Sept. 11.92 UT, five days before the INTEGRAL discovery

(Sunyaev et al., http://atel.caltech.edu/?read=190). The source coor-

dinates are α = 17^{h}54^{m}25^{s}.7, δ = −26^{o}19′58′′ (equinox 2000.0; uncertainty ∼ 10′′), within the reported INTEGRAL error box. The spectrum was consistent with a hydrogen column density of 1.3 × 10^{22} cm^{−2}. The observed flux in the range 2–10 keV was 4.5 × 10^{−12} erg cm^{−2} s^{−1}.

2003 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green